### v2

# Stereo Hand-Object Reconstruction for Human-to-Robot Handover

Abstract
Jointly estimating hand and object shape facilitates the grasping task in human-to-robot handovers. However, relying on hand-crafted prior knowledge about the geometric structure of the object fails when generalising to unseen objects, and depth sensors fail to detect transparent objects such as drinking glasses. In this work, we propose a stereo-based method for hand-object reconstruction that combines single-view reconstructions probabilistically to form a coherent stereo reconstruction. We learn 3D shape priors from a large synthetic hand-object dataset to ensure that our method is generalisable, and use RGB inputs to better capture transparent objects. We show that our method reduces the object Chamfer distance compared to existing RGB based hand-object reconstruction methods on single view and stereo settings. We process the reconstructed hand-object shape with a projection-based outlier removal step and use the output to guide a human-to-robot handover pipeline with wide-baseline stereo RGB cameras. Our hand-object reconstruction enables a robot to successfully receive a diverse range of household objects from the human.
I Introduction
As humans and robots share the same workspace [1], exchanging objects is one of the essential skills to enable smooth collaboration. We consider the scenario where a human hands over various everyday objects to a robot and the robot aims to grasp the object while avoiding the human (human-to-robot handover [2, 3, 4]), and place the object at a predefined delivery location (Fig. 1). Objects can vary in their physical properties, such as appearance, transparency, shape, size, and mass [5]. Container-like objects, such as cups, drinking glasses, or boxes, can be empty or filled with content, affecting the perception of the appearance or the weight of the object [4, 6]. In this highly interactive scenario, hand-object shape reconstruction is an essential task as the object shape informs the robot about how to grasp the object, while the hand shape allows the robot to avoid dangerous contact with the human.
Visual modalities (RGB and depth images) are used in hand-object shape reconstruction for human-to-robot handover. Previous works [7, 3, 8, 9, 10] used depth sensors to capture the pointcloud of the scene and applied segmentation models to extract the pointcloud of the hand and object. However, transparent objects are not easy to be captured by depth sensors [11] and object shapes are incomplete due to partial visibility or occlusions in the scene when using a single camera [3]. Therefore, depth completion methods are devised to fill in the missing depth values [12], but the pointcloud is only filled in on the visible side of the object and a depth camera is still required. As RGB-only cameras are more commonly available than RGB-D cameras, and less expensive, some works use wide-baseline stereo RGB cameras to perform human-to-robot handovers of container-like objects and address partial visibility and occlusions of the object [4, 6]. However, these works are limited to only container-like objects (rotationally symmetric shape), also assuming the object is held upright during the handover due to the presence of a content.
Figure 1: We reconstruct the hand-object pointcloud from stereo RGB input for human-to-robot handover. (1) A safe grasp is selected for the handover and the robot moves in to grasp the object. (2) The object is delivered to a target location on the table. (3) The robot returns to its starting position.
We propose StereoHO, a hand-object reconstruction method for wide-baseline stereo inputs. Our method quantifies the uncertainty of single-view predictions, e.g. caused by self-occlusions, by predicting low-dimensional embeddings of hand and object shapes as a probability distribution over learned codebooks of discrete shape embeddings. Individual predictions from each camera views are aggregated and decoded jointly to form a coherent stereo reconstruction. We train our model on synthetic 3D hand-object data [13] with a large amount of objects and without specific assumptions about the object’s geometry to ensure its generalisation ability. We integrate StereoHO in the first human-to-robot handover pipeline that jointly reconstructs hand and object shapes from stereo RGB inputs while tracking the hand over time. The pipeline also ensures multi-view consistency and reconstruction quality for accurate robot grasp predictions. Experiments [^1] Code and videos of the experiments available at https://qm-ipalab.github.io/StereoHO/ show that StereoHO outperforms previous methods in single-view and stereo settings for object reconstruction, while remaining competitive for hand reconstruction in terms of Chamfer distance. Furthermore, our pipeline can successfully handover objects of diverse shapes and appearances, including filled/empty transparent containers, boxes and thin objects.
II Related work
Recent works in human-to-robot handover  [3, 8, 9, 10, 4, 6, 7] focused on markerless settings that are more applicable to daily human-to-robot interactions. Human grasps were classified into different types to select the corresponding predefined safe robot grasp direction and enable the robot to receive the object [7]. However, this approach does not consider diverse object shapes and limits the grasp types that the robot can handle. Later works improved the perception by capturing the shape of the hand and the object [3, 8, 9, 10, 4, 6], and used existing grasp estimation methods to predict possible grasps on the object shape. Grasps that would cause a collision between the robot and the human are removed for safety. A two-stage system was proposed to perform human-to-robot handovers of everyday objects [3]. The first stage segments the person and the object from the input RGB image. The depth image is lifted into a pointcloud and the background is removed using the segmentation masks. The second stage adds an imaginary plane in the grasping direction to approximate the task as a top-down grasp estimation problem using GG-CNN [14]. Using 6-DoF grasps allows more freedom in the robot’s movement, enabling more natural handovers [8].
Existing methods for human-to-robot handovers [3, 8, 9, 10] and grasp estimation [14, 15], rely on pointcloud as a representation for hand and object shape. However, pointclouds of objects are view-dependent and are incomplete in the presence of occlusions [3] or when the objects are transparent and reflective [11]. Incomplete object shapes or missing pointclouds for transparent objects (Fig. 2) can results in inaccurate grasp estimations. To overcome this issue, depth restoration is a method to complete the pointcloud of transparent objects [11]. However, a depth camera is required which is not as commonly available as RGB cameras. Moreover, the completed pointcloud often only consists of the visible side of the object [12]. An alternative representation for object shapes is Signed Distance Dields (SDFs) that encode into a vector the signed distance between a set of points in 3D and their closest point on the 3D object [16]. As an implicit representation of smooth and continuous surfaces, SDFs distinguish the inside and outside parts of an arbitrary object and are easier to learn compared to pointclouds [16]. None of the existing works has used SDFs for reconstructing object-only or hand-object shapes for more accurate human-to-robot handovers.
Other methods use stereo RGB images as input to reconstruct the shape of the object [4, 6]. The shape of containers, such as drinking glasses and cups, was reconstructed from stereo RGB images by assuming a rotationally symmetric geometry (Fig. 2) and exploiting the predicted segmentation masks of the object [17]. This enables human-to-robot handover by grasping at the centroid of the object [4]. Human safety is considered by estimating the 3D hand keypoints [6]. Robot grasps are sampled along the vertical axis of the object and grasp points that are too close to the hand keypoints are removed. However, these methods do not estimate 6-DoF grasps and thus limit the naturalness of the robot movement as the object must maintain an upright orientation during the handover. The assumption on the object shape is easily invalidated by non-container objects, causing failure in the object shape estimation.
Figure 2: Limitations of existing human-to-robot handover approaches. Depth-based sensing fails on transparent objects (left). Relying on shape priors does not generalise (right).
Joint hand-object reconstruction from RGB images in an interaction scenario is challenging due to severe mutual occlusions between the hand and the object [13]. Large hand-object interaction datasets, both synthetic and collected in real environments [13, 18, 19], have helped improve the reconstruction quality. Existing methods can be grouped into single-view and multi-view, and multi-view methods can be further split into dense and sparse. Single-view reconstruction methods [13, 20, 21, 22] learn 3D shape priors from hand-object datasets, but completing the shape of the unobserved parts of an object remains challenging. Multi-view methods [23, 24, 25, 26, 27] overcome the problems of occlusion and unobserved object parts by increasing the number of input views. Dense multi-view methods [23, 24, 25, 26] use a large number of input views to produce high quality reconstructions but are computationally expensive, limiting the applicability in reactive scenarios. Sparse multi-view methods [27] balances the reconstruction quality of dense methods with the speed of single-view methods. However, training only on synthetic data can affect the model’s ability to generalise to real world scenarios (sim-to-real gap) [27].
Figure 3: Our proposed method, StereoHO, for hand-object reconstruction with two cropped images from a wide-baseline stereo camera. StereoHO predicts the probability distributions, $P_{L}=[P_{HL},P_{OL}]$ and $P_{R}=[P_{HR},P_{OR}]$, over the shape codebooks for each view and combined them into a coherent probability distribution $P=[P_{H},P_{O}]$. The trained SDF decoder transforms $P$ into the hand-object T-SDF. Sampled surface points from the T-SDF (pointcloud $\mathcal{P}$) are projected into each view using the predicted camera projection parameters and outliers are removed by using the segmentation masks to obtain the final pointcloud $\mathcal{P}^{\prime}$. KEYS – $K$: intrinsics calibration parameters, $L$: left view, $R$: right view; $H$: hand, $O$: object.
III Proposed method
In this section, we introduce StereoHO for stereo hand-object reconstruction and our human-to-robot handover pipeline based on StereoHO.
III-A Stereo-based RGB hand-object reconstruction
We design StereoHO to predict the joint surface of the hand ($H$) and object ($O$), represented as pointcloud $\mathcal{P}=[\mathcal{P}_{H},\mathcal{P}_{O}]$, from the RGB images of a stereo-based camera. Images are cropped and centred on the person’s hand holding an object. We formulate stereo-based hand-object shape reconstruction as a view-dependant image-to-shape classification task (encoding) followed by an aggregation step [20] and a transformation of the aggregated shape codes into the final joint hand-object pointcloud (decoding). Fig. 3 illustrates the inference pipeline of StereoHO. To avoid erroneously including background clutter in the reconstructed shape, we include the segmentation masks of the object, $M_{O}$, and the hand, $M_{H}$, as input to the model. We also predict the wrist pose $T_{H}$ for each view to relate the 3D shapes with the cameras views, assuming the camera poses are not available.
For the encoding part, we define a 3D grid centred on the hand wrist $T_{Hv}$, where $v\in\{L,R\}$ is the camera view. For each view $v$, we project the 3D position of each voxel into the image space using the wrist pose and associate each voxel with an image embedding, extracted with a convolutional neural network (e.g., ResNet-18 [28]), at the corresponding pixel location [20]. We use 3D convolutional layers to transform the resulting 4D tensor into a probability distribution $P_{v}\in[0,1]^{D\times D\times D\times C}$ with respect to a codebook of $C$ shape embeddings, $\mathcal{C}=\{e_{c}\in\mathbb{R}^{S}\}_{c=1}^{C}$. We combine the predicted probability distributions from each view into a single coherent prediction via element-wise multiplication: $P=P_{L}P_{R}$. We thus define another 4D tensor of shape embeddings that are selected from $\mathcal{C}$ based on the index of the highest probability for each element of the grid, $E\in\mathbb{R}^{D\times D\times D\times S}$. The same procedure is performed independently for the hand and the object, resulting in the two shape-based tensors, $E_{O}$ and $E_{H}$, respectively.
For the decoding part, we apply hand and object specialised networks, consisting of 3D upconvolutional layers, to obtain the pointclouds $\mathcal{P}_{H}$ and $\mathcal{P}_{O}$ from $E_{O}$ and $E_{H}$, respectively. Each network (SDF decoder) predict SDF values, $\hat{s}\in\mathbb{R}^{D\times D\times D}$. We then sample the Truncated-SDF (T-SDF) at discrete positions and only the values within a threshold of 1 cm to the surface are kept as the predicted pointclouds. We ensure multi-view consistency by re-projecting the predicted hand and object pointcloud $\mathcal{P}_{H}$ and $\mathcal{P}_{O}$ in the images of both camera views using the wrist poses $T_{Hv}$, and by filtering out any outlier $\mathcal{P}_{B}$ with respect to the segmentation masks $M_{H}$ and $M_{O}$, resulting in the final prediction $\mathcal{P}^{\prime}$.
Next, we detail how we learn the codebooks of 3D shape embeddings and the neural network to reconstruct the hand-object pointcloud from the shape embeddings, and the image-to-shape embeddings. As hand-object data is hard to obtain and annotate, we train the components of StereoHO only on synthetic data. Note that the use of segmentation masks facilitate sim-to-real transfer as they are domain-invariant.
Figure 4: Our proposed pipeline for human-to-robot handover. For each frame, hand-object detection extracts bounding boxes $B$ to crop the input images around the hand and then used to segment hand-object masks $M$ and estimate wrist poses $T_{H}$. Our StereoHO uses these outputs along with the image crops to reconstruct the hand-object shape $\mathcal{P}^{\prime}$. We estimate the grasp $g$ on the reconstructed shape and transform it into the robot coordinate system using the wrist poses $T_{HB}$.
III-B Learning discrete 3D shape embeddings
We learn two discrete codebooks, one for the objects $\mathcal{C}_{O}$ and one for the hands $\mathcal{C}_{H}$, by training a 3D Patch-wise Encoding Variational Autoencoder with Vector Quantization [29] on a synthetic hand-object dataset [13]. For each shape (object or hand) in the training set, we transform the shape representation from a 3D mesh to SDFs, discretized in a 3D grid (T-SDFs), $s\in\mathbb{R}^{D\times D\times D}$. The model encodes local patches of the T-SDF with 3D convolutional layers into a shape code $z_{e}\in\mathbb{R}^{S}$, maps each shape code to the closest quantized shape embedding $e\in\mathbb{R}^{S}$ in the codebook (vector quantization [30]), and reconstructs the shape $\hat{s}$. Through mapping the continuous embeddings into a finite set of learned embeddings, vector quantization reduces the high-dimensionality of the shape embeddings [30]. The SDF decoder then retrieves the T-SDF values $\hat{s}$ from $e$.
We use the following loss to train each model by simultaneously learning the autoencoder [^2] We tested the two autoencoders on ObMan [13], achieving a Chamfer distance of 0.015 cm2 and 0.437 cm2 for the hand and object, respectively. parameters and codebook,
$\mathcal{L}_{ae}=|s-\hat{s}|+||sg[z_{e}]-e||^{2}_{2}+\beta||z_{e}-sg[e]||^{2}_{2},$ (1)
where $sg$ is the stop gradient operator. The hyper-parameter $\beta$ controls the scale of the commitment loss (last term), that helps stabilize the training [29]. We initialise the codebook with a uniform distribution. After training the models for object and hand, we only keep the trained SDF decoders and codebooks and link them to the image-to-shape encoder.
III-C Training of the image-to-shape encoder
We train an image-to-shape encoder that transforms the input image and segmentation masks into a probability distributions over the hand and object codebook. We use ResNet-18 [28] pre-trained on ImageNet[31] as our image encoder and 3D convolutional layers as prediction heads for the hand and object. We train all layers with a weighted cross-entropy loss function defined as
$\mathcal{L}_{ce}=-\sum_{c}^{C}{w_{c}}\cdot p_{gt}(c)\cdot\text{log }p(c),$ (2)
where $p_{gt}$ and $w_{c}$ are the ground-truth probability and weight for index $c$. We set a lower weight for the index $c$ corresponding to empty space as the majority of the space surrounding the hand and object in the 3D grid is empty.
III-D Handover with stereo hand-object reconstruction
We integrate StereoHO into a pipeline for human-to-robot handovers (Fig. 4) and assume that i) the handover task starts with the person already holding the object and ii) two cameras are placed on either side of the robot, maximising the visible area of the hand and object.
For each frame of the approaching phase, StereoHO reconstructs the pointclouds of the hand and the object using image crops, pose of the wrist, and segmentation masks from each view. Therefore, for each camera view and for each frame, a hand-object detector [32] estimates the bounding boxes of both the hand and the object, $b_{H},b_{O}\in\mathbb{R}^{4}$. We use Fast Segment Anything [33] to segment the object ($M_{O}$) within the object bounding box [33] and we crop the image around the hand using $b_{H}$. From the image crop, we use FrankMocap [34] to predict the pose of the wrist, $T_{H}$ and the segmentation mask of the hand, $M_{H}$, after extracting the silhouette from the hand shape. We use both segmentation masks to remove the background from the image crop. We triangulate the centroid of the object mask and verify the reprojection error in each view is less than 5 pixels. Note that the pipeline does not perform the handover (failure case) if, within 10 s, no hand-object pair is detected in either of the views or the triangulation is not validated. Since the reconstruction quality can vary across frames, significantly affecting the estimated robot grasp, we compute the Intersection over Union (IoU) of the convex hull of the pointclouds with the segmentation masks, respectively. We update the pointclouds if there is an improvement in reconstruction quality (i.e., IoU $>$ IoU).
To enable the robot to receive the object, we use 6-DoF GraspNet [15] to estimate a set of $N$ candidate grasps, $\mathcal{G}=\{g_{n}\in SE(3)\}_{n=1}^{N}$, on the object pointcloud $\mathcal{P}^{\prime}_{O}$. To ensure human safety, we construct a 3D bounding box of the size of the gripper around each grasp, and we filter out any grasps with points from the hand point cloud $\mathcal{P}^{\prime}_{H}$ in its bounding box. Using the predicted wrist pose in the robot coordinate frame, $T_{HB}=T_{WB}T_{HW}$ where $T_{WB}$ is the robot base transform, we transform the predicted grasps from the hand-wrist to the robot coordinate frame. For each timestep, the robot control selects the grasp $g$ closest to the current gripper 6-DoF pose. We define a standoff position 15 cm away from the grasp point in the direction of the grasp. Once the robot reaches the standoff position, the robot enters a close-loop grasping stage and closes its fingers when the gripper reaches the selected grasp. After grasping, the robot retracts to the standoff position, delivers the object at a predefined location on the table, and returns to its initial position.
III-E Parameter settings
We set the T-SDF dimension to $D=128$. The codebook size is $C=512$ with embedding size $S=128$. For $\mathcal{L}_{ae}$, we set $\beta=1.0$. For $\mathcal{L}_{ce}$, we set $w_{c}=0.25$ for the index $c$ corresponding to empty space and $w_{c}=0.75$ for the rest of the codebook. The robot base transform $T_{WB}$ is obtained using the hand-eye calibration process [35]. We estimate $N=200$ grasp candidates for each handover.
IV Validation
IV-A Datasets
We train StereoHO on the ObMan [13] dataset and test it on DexYCB [18]. The training set of ObMan contains 87,190 images of synthetic handheld objects rendered on top of real images. Objects were selected from ShapeNet [36] based on 8 categories (Fig. 5), such as bottles, bowls, cans, jars, knifes, cellphones, cameras and remote controls, and grasped by the MANO hand model [37]. DexYCB is a multi-view dataset with videos of a person manipulating one of 20 everyday objects selected from YCB-Video [38].
IV-B Handover setup
We follow a benchmarking protocol [4] that includes a setup of the robotic and perception systems, and 288 handover configurations to execute. The setup includes a 6-DoF Universal Robots UR5 robotic arm, equipped with a Robotiq 2F-85 2-finger gripper for grasping the object, and two Intel Realsense D435i cameras mounted on tripods on either side of the robotic arm pointing at the workspace of the robot which is a table with the delivery target marked. The protocol configurations are executed by 4 participants who hand over 3 cups and 1 plastic wine glass, either empty or filled with rice, using 3 different grasp types at 3 handover locations (72 configurations for each participant) [^3] Approval for this research was obtained from the Queen Mary Ethics of Research Committee Ref:QMERC20.525 on 19/10/2021. Consent from each human subject was obtained prior to the experiments.. The protocol instructs to initially place the object on the table at the centre of the delivery target, and then a person picks up and hands the object over to the robot. On the contrary, we assume the object to be already held by the person for our pipeline and therefore we asked participants to hold the object behind a starting line indicated by tape on the wooden table. The robot starts moving towards the object once the human moves the object beyond the starting line. To assess generalisation to different novel objects outside of the benchmark, each participant execute another 24 configurations for 8 household objects at the 3 different handover locations using 1 grasp type (a total of 96 configurations). These additional objects vary in size, topology, and appearance (Fig. 5).
Figure 5: Synthetic object types used for training StereoHO, and objects for testing hand-object reconstruction and handover experiments. For handovers: 3 cups and 1 glass, empty or filled with rice, from a benchmarking protocol [4] (left), and 8 household objects (right) to assess generalisation.
Figure 6: Comparing single-view and stereo hand object reconstructions on DexYCB [18]. For single-view, each reconstruction corresponds to the image on the same row. For stereo, the same reconstruction is shown from two viewpoints. Our method yields less noisy reconstructions and improves both the hand and object shape by combining the single view predictions.
IV-C Performance measures
We use Chamfer distance to evaluate hand and object reconstructions. The Chamfer distance between the ground-truth and predicted point sets $S$ and $\hat{S}$ is
$d_{CD}(S,\hat{S})=\sum_{x\in S}\min_{y\in\hat{S}}||x-y||^{2}_{2}+\sum_{y\in\hat{S}}\min_{x\in S}||x-y||^{2}_{2}$ (3)
To evaluate the performance of the human-to-robot handover methods, we use three measures from the CORSMAL benchmark [4], including delivery location ($\delta$), efficiency ($\gamma$), and delivered mass ($\mu$), and two other measures, including grasping success ($G$) and delivery success ($D$).
For delivery location, we measure the distance from the center of the base of the delivered container to a predefined delivery location on the table as $d$ in mm,
$\delta=\begin{cases}1-\frac{d}{\rho},&\text{if }d<\rho,\\ 0,&\text{otherwise}.\\ \end{cases}$ (4)
We set $\rho=500$ mm as the maximum accepted distance.
We quantify the efficiency of each handover configuration as the amount of time $t$ required from the moment the robot starts moving towards the object to the moment the object is delivered on the table,
$\gamma=\begin{cases}1-\frac{\max(t,\eta)-\eta}{\tau-\eta},&\text{if }t<\tau,\\ 0,&\text{otherwise}.\\ \end{cases}$ (5)
The less time required, the more efficient the method is, and the higher the efficiency score $\gamma$. The score is scaled based on the expected minimum time required $\eta=1$ s and the maximum time allowed $\tau=15$ s.
For the delivered mass, when the container is filled with rice, the robot is expected to deliver the container to the delivery location without spilling the content. We quantify the performance of the delivery based on the difference between the total mass $m$ of the filled container delivered and the total mass $\hat{m}$ of the filled container before the handover,
$\mu=\begin{cases}1-\frac{|m-\hat{m}|}{\hat{m}},&\text{if }|m-\hat{m}|<\hat{m},\\ 0,&\text{otherwise},\\ \end{cases}$ (6)
We measure grasping success as the configurations where the robot is able to grasp and hold the object without dropping it after retracting.
We quantify delivery success of each configuration as the successful placement of the container on the table within the maximum delivery distance $\rho$ with no content spilled:
$D=\begin{cases}1,&\text{if }d<\rho\text{ and }|m-\hat{m}|=0,\\ 0,&\text{otherwise}.\\ \end{cases}$ (7)
Note that since the additional objects are not containers, we do not report on the delivered mass $\mu$ and we remove the condition $|m-\hat{m}|=0$ from the delivery success $D$.
IV-D Hand-object reconstruction from RGB
Since our reconstruction method combines two individual single-view predictions to form the stereo prediction, we evaluate our method in both single-view and stereo settings. We compare StereoHO with IHOI [20] in the single-view setting and with SVHO [27] in the stereo setting. IHOI is an SDF-based method that predicts the shape of the object conditioned on the pose of the hand, and is trained end-to-end. SVHO is a sparse multi-view method that reconstructs hand-object shape and can be easily adapted to use only two views. For a fair comparison, we train our method, IHOI, and SVHO, only on the synthetic data of ObMan, and evaluate the methods using DexYCB [18]. We select the two views closest to the stereo views of our setup and compare the reconstruction quality (Fig. 6). For the single-view setting, we aggregate results from both the left and right views. We separate the testing set into seen and unseen categories based on the similarity in 3D shape between the training and testing objects (Fig. 5).
1010101020202020303030304040404050505050606060607070707011111.51.51.51.522222.52.52.52.53333CD [cm2] (Object)CD [cm2] (Hand)Single seen
10101010202020203030303040404040505050506060606070707070CD [cm2] (Object)Single unseen
1010101020202020303030304040404050505050606060607070707011111.51.51.51.522222.52.52.52.53333CD [cm2] (Object)CD [cm2] (Hand)Stereo seen
10101010202020203030303040404040505050506060606070707070CD [cm2] (Object)Stereo unseen
Figure 7: Single-view and stereo hand-object reconstruction errors on DexYCB [18]. Best results on the bottom-left corner. Legend: CD: chamfer distance,  SVHO [27],  IHOI [20],  Ours (no segmentation mask),  Ours. IHOI, as a single-view reconstruction method, is not evaluated in stereo setting.
TABLE I: Comparison of performance results between hand-object reconstruction methods integrated in a system for human-to-robot handovers of unknown and diverse household objects with varying physical properties (shape, mass, transparency).
Score Method DoF RGB D DC SC HH CORSMAL containers Household objects
Avg C1 C2 C3 C4 Avg O1 O2 O3 O4 O5 O6 O7 O8 Avg C3 C4
$\delta$ CB [6] 3 0.45 0.52 0.58 0.33 0.38 - - - - - - - - - 0.31 0.05 0.57
DB 6 - - - - - 0.06 0.05 0.00 0.05 0.06 0.11 0.00 0.00 0.20 0.01 0.03 0.00
ClearGrasp [11] 6 - - - - - 0.01 0.00 0.00 0.00 0.00 0.00 0.05 0.00 0.08 0.00 0.00 0.00
StereoHO 6 0.48 0.26 0.65 0.59 0.42 0.18 0.18 0.11 0.06 0.25 0.31 0.09 0.18 0.23 0.55 0.65 0.45
$\gamma$ CB [6] 3 0.29 0.31 0.35 0.26 0.25 - - - - - - - - - 0.20 0.16 0.24
DB 6 - - - - - 0.29 0.34 0.27 0.30 0.27 0.37 0.27 0.10 0.38 0.13 0.10 0.16
ClearGrasp [11] 6 - - - - - 0.14 0.15 0.05 0.06 0.14 0.15 0.28 0.14 0.16 0.09 0.10 0.08
StereoHO 6 0.10 0.12 0.09 0.10 0.09 0.21 0.09 0.07 0.07 0.14 0.13 0.39 0.44 0.35 0.09 0.09 0.09
$\mu$ CB [6] 3 0.78 0.77 0.81 0.77 0.78 - - - - - - - - - - - -
StereoHO 6 0.74 0.53 0.87 0.84 0.73 - - - - - - - - - - - -
$G$ CB [6] 3 0.79 0.77 0.81 0.79 0.80 - - - - - - - - - 0.83 0.66 1.00
DB 6 - - - - - 0.41 0.16 0.16 0.50 0.50 0.58 0.50 0.08 0.83 0.08 0.08 0.08
ClearGrasp [11] 6 - - - - - 0.15 0.00 0.00 0.08 0.25 0.16 0.41 0.00 0.33 0.20 0.33 0.08
StereoHO 6 0.75 0.55 0.88 0.84 0.73 0.75 0.50 0.66 0.75 0.75 0.83 0.75 1.00 0.75 0.91 1.00 0.83
$D$ CB [6] 3 0.67 0.73 0.77 0.50 0.68 - - - - - - - - - 0.50 0.08 0.91
DB 6 - - - - - 0.20 0.08 0.16 0.33 0.25 0.25 0.00 0.00 0.58 0.04 0.08 0.00
ClearGrasp [11] 6 - - - - - 0.05 0.00 0.00 0.00 0.00 0.00 0.16 0.00 0.25 0.00 0.00 0.00
StereoHO 6 0.66 0.36 0.86 0.84 0.59 0.50 0.41 0.58 0.50 0.58 0.83 0.08 0.58 0.41 0.83 1.00 0.66
Results of our proposed method highlighted in gray. Objects highlighted based on their transparency (transparent, opaque, mixed). Best scores highlighted in bold for each column and each performance measure. KEY – DoF: degrees of freedom, RGB: colour input, D: depth input, DC: depth completion, SC: stereo camera, HH: handheld object assumption, Avg: average, C1: small white cup, C2: red cup, C3: beer cup, C4: wine glass, O1: tea box, O2: cracker box, O3: tape, O4: screwdriver, O5: CD, O6: cloth, O7: cap, O8: spray bottle; $\delta$: delivery location, $\gamma$: efficiency, $\mu$: delivered mass, $G$: grasping success, $D$: delivery success, : not used, : used; -: not applicable.
In the single-view setting (Fig. 7), our method outperforms IHOI on object reconstruction by 9.71 cm2 and 12.52 cm2, while slightly underperforming on hand reconstruction by 1.10 cm2 and 1.21 cm2 in the seen and unseen categories, respectively. In the stereo setting, our method outperforms SVHO on object reconstruction by 35.43 cm2 and 32.27 cm2, while performing slightly better on the hand reconstruction by 0.51 cm2 and 0.13 cm2 in the seen and unseen categories, respectively. Compared with a version of our method without segmentation masks as input, our method improves the object reconstruction while sacrificing a small amount of performance in hand reconstruction. This also results in less noisy reconstructions as shown in Fig. 6 when comparing with IHOI in the single-view setting.
IV-E Human-to-robot handovers
We compare StereoHO with a depth-based baseline (DB), ClearGrasp [11], and the CORSMAL baseline (CB) [6]. DB uses the pointcloud observed from the left camera. ClearGrasp is a depth completion method to handle transparent objects using depth-based sensing. For fairness, DB and ClearGrasp are integrated in the same handover pipeline. CB [6] is a stereo-based pipeline specifically designed for handovers of rotationally symmetric container objects. Unlike our pipeline, CB assumes the object is initially placed on the table, following the original protocol [4], and is upright throughout the entire handover process. The object is detected using Mask R-CNN [39], pretrained on COCO [40], and the object shape is estimated with a multi-view iterative fitting algorithm given a cylindrical prior [17]. For each frame, hand keypoints are estimated using OpenPose [41] and the 3D object centroid is triangulated from the 2D centroid in the left and right views after tracking the object mask in each view with SiamMask [42]. Grasp points are generated along the height of the container and those ones close to the hand keypoints on the vertical axis are removed. The robot starts moving once the object is lifted off the table.
Figure 8: Comparing hand (red) and object (green) reconstructions, overlaid on the pointcloud obtained by the camera.
Tab. I compares the handover performance of the methods. For the container objects from the CORSMAL benchmark [4], both CB and our method achieve over 75% grasping success and 65% delivery success. This shows that assuming a cylindrical prior and a mostly upright pose throughout the task is reasonable when handing over cups and drinking glasses [6]. When handing over general household objects, our method achieved 75% grasping success on the opaque objects and 91% grasping success on the transparent containers, significantly outperforming DB and ClearGrasp. Comparing the hand and object shape reconstructed by each method (Fig. 8), DB is easily affected by transparency of the object and noisy depth measurements, leaving holes in the reconstructed shape. ClearGrasp can fill in the missing depth value but the estimated value is often inaccurate and noisy. Our method produces a complete shape which helps the grasp estimation method to produce successful grasp candidates for the handover.
Figure 9: Examples of successful human-to-robot handovers of household objects using StereoHO, and failure cases due to grasping failure or spilled filling.
Overall, we showed that our method can perform human-to-robot handovers on general household objects (Fig. 9), outperforming depth-based methods and performing on par on containers with a method designed specifically for the handover of containers. Specifically, our method can successfully receive objects with challenging shapes, such as thin (screwdriver), deformable (cloth), and reflective objects (CD), whereas previous stereo-based method fail to generalise to non-container objects, and depth-based methods fail to receive transparent objects. Failures occur when the reconstruction of the object shape is inaccurate. For example, the robot often mistakenly try to grasp the cracker box from the wider side (Fig. 9), which is larger than the gripper width. Failure can also occur when the robot loses track of the hand due to occlusions by the object or the robot arm itself.
V Conclusion
In this paper, we proposed a joint hand-object reconstruction method to enable human-to-robot handovers using wide baseline stereo RGB cameras. Our reconstruction method outperforms existing methods in both single-view and stereo settings. Handover experiments showed that our method enables a robot to receive empty and filled containers, and can generalize to other household objects. Future work includes improving the speed of handovers and the quality of hand-object reconstructions.
ACKNOWLEDGMENT
This project made use of time on Tier 2 HPC facility JADE2, funded by EPSRC (EP/T022205/1).